OpenStudy (anonymous):

Help need quick. Evaluate - 5+|1-4|(3-6)²÷3 A. 6 B. 11 C. 14 D. 24

OpenStudy (anonymous):

abs(1-4)=abs(-3)=3 (3-6)^2=(-3)^2=9

OpenStudy (anonymous):

Now, use PEDMAS

OpenStudy (yttrium):

B is the answer for you checking. :)